SUTURED MANIFOLD HIERARCHIES, ESSENTIAL LAMINATIONS, AND DEHN SURGERY

YING-QING WU

Abstract. We use sutured manifold theory, essential laminations and essential branched surfaces to establish the upper bounds of distances between certain types of nonsimple Dehn surgery slopes. This is the revised version of an earlier preprint Dehn surgery and simple manifolds.